• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name ellie

    Log entry time 15:59:50 on October 07, 2009

    Entry number 293687

    keyword=Day Shift Summary

    10/06/09 Day Shift Summary

    On shift:
    SL: Ellie Long
    TO: Mustafa Canan
    3rd: Maria Leda Sperduto

    · 08:00 Good morning, Hall A! Starting shift in HAPPEX Production mode. It's nice when things start out working. IHWP=OUT, Slug #19

    · 09:55 Changed beam current to 15uA for Compton work.

    · 10:22 Changed beam current to 100uA. Starting Production Runs.

    · 10:50 LeCroy HV Alarm sounded from Card 11, Channel 1. Turned off and back on. Problem fixed.

    · 12:35 Compton ROC 6 stopped responding. Attempted reboot, failed. Tried again, failed. Called RC and expert to look at it. Running HAPPEX production but not Compton.

    · 13:15 Called MCC to go to Controlled Access to move the Compton Detector closer to the beam.

    · 13:48 Compton work done, going to Beam Permit.

    · 16:00 Ending shift in Production Mode. Good luck, night owls!

    Production Runs:
     Run #  Beam Current  # of Good Pairs  Cummulative # of Good Pairs  Comments